The results of a history exam reveals a normal distribution of marks with an average of 67% and a standard deviation of 5%. READ THE QUESTIONS CAREFULLY.


The TOP 94.52% of students scored at least what grade on the test?

Answer for part 1

%

The bottom 0.02% of students scored equal to or lower than what mark?

Answer for part 2

%


1
Expert's answer
2021-04-20T07:32:45-0400

"\mu=67,\sigma=5"

n=100

TOP 94.52% of students scored at least marks= "\mu-z_{0.0025}\dfrac{\sigma}{\sqrt{n}}"


"=67-1.96\dfrac{5}{\sqrt{100}}=67-0.98=66.02" %


The bottom 0.02% of students scored equal to or lower than what mark

"=\mu+z_{0.0001}\dfrac{\sigma}{\sqrt{n}}"


"=67+3.042\times \dfrac{5}{\sqrt{100}}=67+1.521=68.521" %